Problem 3-14 Schedule of Cost of Goods Manufactured; Overhead Analysis [LO3-3, LO3-4] Gitano Products operates a job-order costing system and applies overhead cost to jobs on the basis of direct materials used in production (not on the basis of raw materials purchased). Its predetermined overhead rate was based on a cost formula that estimated $128,800 of manufacturing overhead for an estimated allocation base of $92,000 direct material dollars to be used in production. The company has provided the following data for the just completed year: Purchase of raw materials $ 136,000 Direct labor cost $ 88,000 Manufacturing overhead costs: Indirect labor $ 117,900 Property taxes $ 8,200 Depreciation of equipment $ 19,000 Maintenance $ 13,000 Insurance $ 8,400 Rent, building $ 40,000 Beginning Ending Raw Materials $ 26,000 $ 17,000 Work in Process $ 48,000 $ 36,000 Finished Goods $ 74,000 $ 61,000 Required: 1. Compute the predetermined overhead rate for the year. 2. Compute the amount of underapplied or overapplied overhead for the year. 3. Prepare a schedule of cost of goods manufactured for the year. Assume all raw materials are used in production as direct materials. 4. Compute the unadjusted cost of goods sold for the year. Do not include any underapplied or overapplied overhead in your answer. 5. Assume that the $36,000 ending balance in Work in Process includes $8,000 of direct materials. Given this assumption, supply the information missing below:
Answer #1

Required: 1. Compute the predetermined overhead rate for the year.

a) Predetermine overhead rate = 128800/92000 = 1.4 per dollar direct material used

2) Direct material used = 26000+136000-17000 = 145000

Overhead applied = 145000*1.4 = 203000

Actual overhead = 117900+8200+19000+13000+8400+40000 = 206500

Under applied overhead = 203000-206500 = 3500

3) Schedule of cost of goods sold

Direct material used
Beginning raw material 26000
Raw material purchase 136000
Less: Ending raw material -17000
Direct material used 145000
Direct labor 88000
Overhead applied 203000
Total manufacturing cost 436000
Beginning WIP 48000
Total Cost of work in process 484000
Less: Ending WIP -36000
Cost of goods manufactured 448000

Schedule of cost of goods sold

Beginning finished goods 74000
Cost of goods manufactured 448000
Cost of goods available for sale 522000
Less: Ending finished goods -61000
Unadjusted cost of goods sold 461000

Ending balance in WIP = 36000

Direct material = 8000

Overhead applied = 8000*1.4 = 11200

Direct labor = (36000-8000-11200) = 16800